OpenStudy (anonymous):

REALLY NEED HELP WITH THIS ESSAY QUESTION!!! WILL MEDAL AND FAN PLEASE HELP The elevation of city A is 72 feet above sea level; the elevation of city C is 24 feet above sea level. The elevation of city D is 1 over 4 of the elevation of city C. Part A: Write an expression to represent the elevation, in feet, of city D. (5 points) Part B: Simplify the expression and explain why the answer is negative or positive. (5 points)

OpenStudy (anonymous):

@imqwerty @Jadedry

OpenStudy (rccb):

Part A : D = (1/4) * C D = 1/4 * 24 D = 24/4 = 6 ft above sea level Part B : Your answer is positive because D is still above sea level...if it would have been below sea level, then it would have been negative

OpenStudy (anonymous):

What does the * mean?

OpenStudy (rccb):

multiplication, so in words, 1 over 4 times c